find the vertex the line of symmetry the maximum and minimum value of the quadratic function. F(x)=x^2-12x-3
Answer by Alan3354(69443)

The line of symmetry is x = -b/2a
x = 12/2
x = 6
------
The minimum is F(6) = 36 - 72 - 3
= -39